1303 +3.00 C -2.00 nC 0.0 30 +3.00 nC N/C direction
Answer #1

Since the charges and their distance are equal so the vertical component of electric field due to two 3nC charge get balanced, hence the net electric field is given only along x axis

E = 2* cosine component of electric field due to 3 NC - electric field due to (-2 nC)

E = 2 * cos 30* k q1 / r^2 - k q2 / r^2

E = 9*10^9* 10^-9 / 0.0364^2 * ( 2* cos 30* 3 - 2)

E = 21710.37 N/C

Directed along +x axis

=======

Force  

F = qE = 5.4*10^-9* 21710.37 = 1.172*10^-4 N/C

Directed along - x axis

======

Comment in case any doubt please.. Goodluck